# Multilingual European Language Models: Benchmarking Approaches and Challenges

### v1

@@ 1 Introduction
Recent advances in large language models (LLMs) have demonstrated very strong capabilities across a wide range of natural language processing (NLP) tasks. The performance of these models is measured and quantified using well-established benchmarks in which the models have to perform specific tasks such as, among others, reasoning, question answering (QA), or common-sense inference on English text (see, for example, Hendrycks et al., 2021; Lin et al., 2022; Zellers et al., 2019). The benchmarks consist of text labeled by human annotators Ding et al. (2024) or they were developed with the help of machine learning techniques that annotate or label the original text data automatically Pfister and Hotho (2024).
@@
Recently, researchers have been using multilingual benchmarks that are comparable and quantifiable across languages Touvron et al. (2023). For instance, in late 2024, three multilingual LLMs were published that cover all 24 European languages Ali et al. (2024); Martins et al. (2024); Gonzalez-Agirre et al. (2025). These LLMs were evaluated using English (MMLU, Hellaswag, TruthfulQA) and a few multilingual benchmarks Hendrycks et al. (2021); Zellers et al. (2019); Lin et al. (2022); Bandarkar et al. (2024). These evaluations were performed using machine-translated versions of well-established English benchmarks regarding multilingual reasoning, QA, or common-sense inference Thellmann et al. (2024). The translated benchmarks are, like the original English versions, multiple-choice problems, where the model has to choose the correct answer or output from four possible answer options. The correct answer can be calculated by generating the correct output Zellers et al. (2019) or measuring the log-likelihood of each possible answer Wiland et al. (2024). The benefit of automatically translated benchmarks is that, besides cost- and time-efficiency in terms of dataset generation, the scores are comparable across languages. [^1] https://huggingface.co/spaces/occiglot/euro-llm-leaderboard
@@ 3.3 Challenge 3: Cultural Bias
Most of the benchmarks evaluate the reasoning capabilities of LLMs. However, reasoning, behavior, and communication are shaped by culture Tao et al. (2024). Therefore, cultural biases in multilingual datasets present substantial obstacles with regard to their scope, validity and reliability as global benchmarks Singh et al. (2024). Cultural biases will not be mitigated if benchmarks are automatically translated from English into other languages. The published benchmarks apply two different options to address this limitation. First, human annotators evaluate cultural biases present in the original dataset. Global-MMLU, for instance, improved the quality of a multilingual MMLU by engaging with professional and community annotators that label samples as culturally-sensitive or culturally-agnostic Singh et al. (2024). Using annotators to verify translations and evaluate cultural bias supports the effectiveness of a multilingual benchmark.
@@
The integration of diverse and culturally grounded knowledge is necessary to achieve a certain level of inclusivity and fairness for multilingual evaluations Singh et al. (2024). Mitigating cultural bias from benchmarks is fundamental for achieving a certain level of reliability for global benchmarks, but – as already mentioned – culture shapes reasoning and is, therefore, an essential component in and for communication and for solving multilingual tasks.

### v2

@@ 1 Introduction
Recent advances in large language models (LLMs) have demonstrated very strong capabilities across a wide range of natural language processing (NLP) tasks. Their performance is measured and quantified using well-established benchmarks in which the models have to perform tasks such as, among others, reasoning, question answering (QA), or common-sense inference on English text (see, e. g., Hendrycks et al., 2021; Lin et al., 2022; Zellers et al., 2019). The benchmarks consist of text labeled by human annotators Ding et al. (2024) or they were developed with the help of machine learning techniques that annotate or label the original text data automatically Pfister and Hotho (2024).
@@
Recently, researchers have been using multilingual benchmarks that are comparable across languages Touvron et al. (2023). For instance, in late 2024, three multilingual LLMs were published that cover all 24 European languages Ali et al. (2024); Martins et al. (2024); Gonzalez-Agirre et al. (2025). EuroLLM and Teuken-7B were evaluated using English (MMLU, Hellaswag, TruthfulQA) and a few multilingual benchmarks Hendrycks et al. (2021); Zellers et al. (2019); Lin et al. (2022); Bandarkar et al. (2024). These evaluations were performed using machine-translated versions of English benchmarks regarding multilingual reasoning, QA, or common-sense inference Thellmann et al. (2024). Salamandra, however, was evaluated on multilingual datasets that had been human-annotated or human-translated, but do not cover all European languages Baucells et al. (2025).
The translated benchmarks are, like the original English versions, multiple-choice problems, where the model has to choose the correct answer or output from four possible answer options. The correct answer can be calculated by generating the correct output Zellers et al. (2019) or measuring the log-likelihood of each possible answer Wiland et al. (2024). The benefit of automatically translated benchmarks is that, besides cost- and time-efficiency in terms of dataset generation, the scores are comparable across languages. [^1] https://huggingface.co/spaces/occiglot/euro-llm-leaderboard
@@ 3.3 Challenge 3: Cultural Bias
Most benchmarks evaluate the reasoning capabilities of LLMs. However, reasoning, behavior, and communication are shaped by culture Tao et al. (2024). Therefore, cultural biases in multilingual datasets present substantial obstacles with regard to their scope, validity and reliability as global benchmarks Singh et al. (2024). Cultural biases will not be mitigated if benchmarks are automatically translated from English into other languages. The published benchmarks apply two different options to address this limitation. First, human annotators evaluate cultural biases present in the original dataset. Global-MMLU, for instance, improved the quality of a multilingual MMLU by engaging with professional and community annotators that label samples as culturally-sensitive or culturally-agnostic Singh et al. (2024). Using annotators to verify translations and evaluate cultural bias supports the effectiveness of a multilingual benchmark.
@@
The integration of diverse and culturally grounded knowledge is necessary to achieve a certain level of inclusivity and fairness for multilingual evaluations Singh et al. (2024). Mitigating cultural bias from benchmarks is fundamental for achieving a certain level of reliability for global benchmarks, but – as mentioned – culture shapes reasoning and is, therefore, an essential component in and for communication and for solving multilingual tasks.